# PAPER: Privacy-Preserving ResNet Models using Low-Degree Polynomial Approximations and Structural Optimizations on Leveled FHE

### v1

@@ 2.2 Training Strategy
Quantization-Aware Polynomial Fitting. To avoid coefficient truncation, we integrate fixed-point constraints directly into coefficient estimation using quantization-aware polynomial fitting (Diaa et al., 2024). Given an approximation interval [c,c][-c,c] and fixed-point precision with $b$ fractional bits, we define a quantized input domain for polynomial fitting as: $X=\{x\in[-c,c]\mid x=k\cdot 2^{-b},\ k\in\mathbb{Z}\}.$ For each input $x_{i}\in X$, we compute scaled ReLU outputs $Y_{i}=2^{b}\cdot ReLU(x_{i})$ so that regression targets are integers, which reduces risk of precision loss during optimization. We build a matrix $B\in\mathbb{R}^{M\times(d+1)}$ with entries $B_{i,k}=x_{i}^{k}\ \text{for }0\leq k\leq d.$ The coefficients $A=[a_{0},\ldots,a_{d}]^{\top}\in\mathbb{Z}^{d+1}$ are obtained by solving bounded integer least-squares problem: $\min_{A\in\mathbb{Z}^{d+1}}\|BA-Y\|_{2}^{2}\ \text{subject to }a_{k}\in[-2^{b-1},2^{b-1}-1].$ The resulting fixed-point polynomial is $p_{d}(x)=\sum_{k=0}^{d}(\frac{a_{k}}{2^{b}})x^{k}.$
@@
$\mathcal{L}_{\mathcal{B}}\;=\;\;\underbrace{\frac{1}{|\mathcal{B}|}\sum_{(x,y)\in\mathcal{B}}\ell_{\mathrm{CE}}\left(g\left(x\right),y\right)}_{\text{classification loss}}\;+\;\zeta\;\underbrace{\frac{1}{L}\sum_{l=1}^{L}\left\|z_{p}^{(l)}\;-\;\mathrm{clip}\left(z_{p}^{(l)};[-c,c]\right)\right\|_{2}}_{\text{clip-range penalty}}$ (1)
where $\ell_{\mathrm{CE}}$ is the cross-entropy loss, $z_{p}^{(l)}$ are pre-activations at layer $l$, i.e., outputs of affine transformations $W^{(l)}h^{(l-1)}_{p_{d}}(x)$ before the polynomial activation. The function $\mathrm{clip}(z;[-c,c])$ is applied element-wise and defined as: $\mathrm{clip}(z;[-c,c])_{i}=\max(-c,\min(z_{i},c))$. The second loss term penalizes pre-activations that lie outside the interval [c,c][-c,c], with strength controlled by the regularization parameter $\zeta>0$. See §C for a proof of the penalty function’s correctness. Training can still become unstable for two reasons: (i) in early epochs pre-activations can grow unbounded before the model learns to contain them and (ii) using full regularization weight $\zeta$ from the start can let the penalty dominate and destabilize optimization. As additional strategies, we also consider clipping pre-activations during training and introducing a warm-up schedule for $\zeta$ (see §D for more details).
@@ 3.1 Node Fusing
Our first optimization is node fusing, which merges consecutive computational nodes into a single equivalent function. This reduces both number of operations and multiplicative depth. Formally, two sequential nodes $F(G(x))$ are replaced with $H(x)$ such that $F(G(x))=H(x)$, with $H(x)$ more efficient to evaluate. Since our method uses batch normalization, recall its polynomial form $B(x)=b_{1}x+b_{0},\;b_{1}=\frac{\gamma}{\sigma},\;b_{0}=\beta_{b}-b_{1}\mu$, where $\mu,\sigma$ are the input mean and standard deviation, and $\gamma,\beta_{b}$ are learnable parameters. We outline four common fusion cases. See §E.1 for full derivations.
@@
Case 2: $B(C(x))\mapsto C(x)$. When batch normalization follows a convolution $C(\cdot)$, the two can be fused (Markuš, 2018). A convolution is given by $C(x)=\sum_{i}w_{i}x_{i}+\beta_{c}$, where $w_{i}$ are the convolution weights and $\beta_{c}$ is the bias. The fused form becomes a single convolution $B(C(x))=\sum_{i}\omega_{i}x_{i}+\alpha$, with fused coefficients $\omega_{i}=b_{1}w_{i},\;\alpha=b_{1}\beta_{c}+b_{0}$.
Case 3: $P(B_{X}(x)+B_{Y}(y))\mapsto S(x,y)$. In residual networks, skip connections often merge two batch-normalized branches by summing them before applying an activation $P(\cdot)$. Specifically, $P(B_{X}(x)+B_{Y}(y))$. Here, $B_{X}$ and $B_{Y}$ are independent batch normalization layers to input $x$ and $y$. This structure can be fused into a single quadratic bivariate polynomial, which we call polyskip: $S(x,y)=d_{X2}x^{2}+d_{Y2}y^{2}+d_{XY}xy+d_{X}x+d_{Y}y+d_{0}$, with coefficients $d_{X2}=c_{2}b_{X1}^{2},\;d_{Y2}=c_{2}b_{Y1}^{2},\;d_{XY}=2c_{2}b_{X1}b_{Y1}$, $d_{X}=b_{X1}(2c_{2}(b_{X0}+b_{Y0})+c_{1}),\;d_{Y}=b_{Y1}(2c_{2}(b_{X0}+b_{Y0})+c_{1})$, $d_{0}=c_{2}(b_{X0}+b_{Y0})^{2}+c_{1}(b_{X0}+b_{Y0})+c_{0}$.
Case 4: $P(B_{X}(x)+y)\mapsto S(x,y)$. Some skip connections use identity shortcuts, where the input $y$ is added directly to the batch-normalized branch $B_{X}(x)$ before applying the activation. This structure can be fused into a quadratic bivariate polynomial: $S(x,y)=d_{X2}x^{2}+d_{Y2}y^{2}+d_{XY}xy+d_{X}x+d_{Y}y+d_{0}$,with coefficients $d_{X2}=c_{2}b_{X1}^{2},\;d_{Y2}=c_{2},\;d_{XY}=2c_{2}b_{X1}$, $d_{X}=b_{X1}(2c_{2}b_{X0}+c_{1}),\;d_{Y}=2c_{2}b_{X0}+c_{1}$, $d_{0}=c_{2}b_{X0}^{2}+c_{1}b_{X0}+c_{0}$.
Node fusing collapses batch normalization, activations, convolutions, and skip connections into single fused polynomials, eliminating redundant nodes and reducing multiplicative depth.
@@ 3.2 Weight Redistribution
Our second optimization is weight redistribution, a technique that redistributes weights across the network while maintaining functional equivalence. The goal is to reduce the multiplicative depth of certain functions by one. It specifically targets average pooling, batch normalization, and polynomial activations. For average pooling, redistribution eliminates the normalization step by setting the divisor to one. For polynomial functions, the highest-order coefficient is set to one. For example, the polynomial activation $c_{2}x^{2}+c_{1}x+c_{0}$ is transformed into $x^{2}+p_{1}x+p_{0}$, reducing its multiplicative depth from two to one, the theoretical minimum for an activation function. These transformations alone break model equivalence. To maintain it, other nodes in the network must compensate for the change. We refer to the nodes initiating weight redistribution as donors, and the nodes adjusting their parameters to preserve equivalence as receivers. We represent the network as a directed graph and traverse it, identifying all eligible donor nodes. For each donor, redistribution can be applied either forward or backward, affecting receivers among the donor’s successors or predecessors, respectively.
@@
Update Forward. Donors. We first consider donor updates in the forward direction. Our goal is to construct a normalized function $\bar{F}(x)$ such that $\upsilon\bar{F}(x)=F(x)$, where $F(x)$ is the original donor function and $\upsilon$ is the update term to be propagated to receivers. The average pooling operation is given by $\mu(x)=k^{-1}\sum_{i}{x_{i}}$, where $k$ is the kernel size. In the normalized function we set $k=1$, reducing the operation to $\bar{\mu}(x)=\sum_{i}x_{i}$. Thus, for the equality to be valid, we have $\upsilon=k^{-1}$.
@@
Batch normalization and polynomial activation are polynomial functions. For a univariate polynomial $P(x)=\sum_{i=0}^{d}{c_{i}x^{i}}$ of degree $d$, and a normalized polynomial $\bar{P}(x)=x^{d}+\sum_{i=0}^{d-1}{\bar{c}_{i}x^{i}}$, where $\bar{c}_{i}$ are the normalized coefficients, we have $\upsilon=c_{d}$ and $\bar{c}_{i}=c_{i}\upsilon^{-1}$ for $\upsilon\bar{P}(x)=P(x)$ to hold. The extension to the bivariate polynomial activation is analogous and deferred to §E.2.
@@
Receivers. After a donor update, receivers must be adjusted to maintain model equivalence. The original composed function is $G(F(\cdot))$, where $F(\cdot)$ is the donor and $G(\cdot)$ is the receiver. Updating the donor to $\bar{F}(\cdot)$ without compensating the receiver results in $G(\upsilon^{-1}F(\cdot))$, breaking model equivalence. To restore it, we modify the receiver so that $\bar{G}(\bar{F}(\cdot))=\bar{G}(\upsilon^{-1}F(\cdot))=G(F(\cdot))$.
@@
Receivers fall into two categories: kernel functions and polynomial functions. Kernel functions share the form $K(x)=\sum_{i}w_{i}x_{i}+\beta$. To maintain model equivalence, we must update the kernel function such that $\bar{K}(\upsilon^{-1}x)=K(x)$. For that, we must set $\bar{w}_{i}=w_{i}\upsilon$ and $\bar{\beta}=\beta$.
@@
Polynomial functions update as $\bar{P}(\upsilon^{-1}x)=P(x)$, which implies that $\bar{c}_{i}=c_{i}\upsilon^{i}$. For the bivariate case, in $\bar{S}(\upsilon^{-1}x,y)=S(x,y)$ the index $i$ corresponds to the exponent of $x$, while in $\bar{S}(x,\upsilon^{-1}y)=S(x,y)$ it corresponds to the exponent of $y$.
@@
Update Backward. Donors. Backward updates are essentially the inverse of forward updates. For average pooling, the update is identical, since $\bar{\mu}(\upsilon x)=\upsilon\bar{\mu}(x)$, while for polynomial donors, it must hold that $\bar{P}(\upsilon x)=P(x)$; hence, $\upsilon=c_{d}^{1/d}$ and $\bar{c}_{i}=c_{i}\upsilon^{-i}$.
@@
Receivers. Receivers must satisfy $\bar{G}(x)=\upsilon G(x)$ to preserve model equivalence. Kernel receivers are updated as $\bar{K}(x)=\upsilon K(x)$, which implies that $\bar{w}_{i}=w_{i}\upsilon$ and $\bar{\beta}=\beta\upsilon$. And polynomial receivers update according to $\bar{P}(x)=\upsilon P(x)$; therefore, $\bar{c}_{i}=c_{i}\upsilon$.

### v2

@@ 3.2 Training Strategy
Quantization-Aware Polynomial Fitting. To avoid precision loss due to coefficient truncation, we incorporate fixed-point constraints directly into coefficient estimation procedure using quantization-aware polynomial fitting [11]. Given an approximation interval [c,c][-c,c] and fixed-point precision with $b$ fractional bits, we define a quantized input domain for polynomial fitting as: $X=\{x\in[-c,c]\mid x=k\cdot 2^{-b},\ k\in\mathbb{Z}\}.$ For each input $x_{i}\in X$, we compute scaled ReLU outputs $Y_{i}=2^{b}\cdot ReLU(x_{i})$ so that regression targets are integers, which reduces risk of precision loss during optimization. We construct a matrix $B\in\mathbb{R}^{M\times(d+1)}$ with entries $B_{i,k}=x_{i}^{k}\ \text{for }0\leq k\leq d.$ The coefficients $A=[a_{0},\ldots,a_{d}]^{\top}\in\mathbb{Z}^{d+1}$ are obtained by solving bounded integer least-squares problem: $\min_{A\in\mathbb{Z}^{d+1}}\|BA-Y\|_{2}^{2}\ \text{subject to }a_{k}\in[-2^{b-1},2^{b-1}-1].$ The resulting fixed-point polynomial is $p_{d}(x)=\sum_{k=0}^{d}(a_{k}/2^{b})x^{k}.$
@@
$\begin{split}\mathcal{L}_{\mathcal{B}}&=\underbrace{\frac{1}{|\mathcal{B}|}\sum_{(x,y)\in\mathcal{B}}\ell_{\mathrm{CE}}\!\left(g(x),y\right)}_{\text{classification loss}}\\ &\quad+\;\zeta\;\underbrace{\frac{1}{L}\sum_{l=1}^{L}\left\|z_{p}^{(l)}-\mathrm{clip}\!\left(z_{p}^{(l)};[-c,c]\right)\right\|_{2}}_{\text{clip-range penalty}}\end{split}$ (1)
where $\ell_{\mathrm{CE}}$ is the cross-entropy loss, $z_{p}^{(l)}$ are pre-activations at layer $l$, i.e., outputs of affine transformations $W^{(l)}h^{(l-1)}_{p_{d}}(x)$ before the polynomial activation. The function $\mathrm{clip}(z;[-c,c])$ is applied element-wise and defined as: $\mathrm{clip}(z;[-c,c])_{i}=\max(-c,\min(z_{i},c))$. The second loss term penalizes pre-activations that lie outside [c,c][-c,c], with strength controlled by the regularization parameter $\zeta>0$. See §A for a proof of the penalty function’s correctness. Training can still become unstable for two reasons: (i) in early epochs pre-activations can grow unbounded before the model learns to contain them and (ii) using full regularization weight $\zeta$ from the start can let the penalty dominate and destabilize optimization. As additional strategies, we also consider clipping pre-activations during training and introducing a warm-up schedule for $\zeta$.
Pre-Activation Clipping. During early training stages, network weights remain close to their random initialization, and the optimization process does not immediately constrain the pre-activation values to lie within the target approximation interval [c,c][-c,c]. As a result, unbounded pre-activations may arise before the model learns to keep them within range, causing training instability that can degrade model behavior. To prevent this, we restrict pre-activations to the interval [c,c][-c,c] before evaluating the polynomial activation, which is achieved through a clipping strategy $\mathrm{clip}(z;[-c,c])$ [11]:
Importantly, this clipping operation is applied only during training to stabilize learning. It is performed after computing the clip-range penalty to ensure that gradients from the regularization term, which encourages the network to keep pre-activations within the approximation interval [c,c][-c,c], are preserved and not masked by the clipping. At inference time, the clipping function is removed: $h^{(l)}_{p_{d}}(x)\;=\;p_{d}\left(W^{(l)}h^{(l-1)}_{p_{d}}(x)\right)$. The model, having learned to constrain pre-activations during training, is expected to remain within the approximation interval without explicit clipping at inference.
Regularization Warm-up. While activation regularization and pre-activation clipping are both necessary for training stability, applying the full regularization strength from the outset can lead to numerical instability, particularly in larger models. In the early epochs, many pre-activation values lie outside the target interval [c,c][-c,c] across several layers. The clip-range penalty adds a contribution from every such layer, so the penalty term becomes very large. In extreme cases, this can cause the total loss to become numerically undefined.
To address this issue, the regularization strength $\zeta$ is progressively increased over the initial epochs, which is implemented through a regularization warm up schedule [11]. Let $T_{warm}$ denote the total number of warm-up epochs. For each epoch $t$, we define a time-dependent regularization weight $\zeta_{t}$:
where $\{\alpha_{t}\}_{t=1}^{T_{warm}}$ is a fixed sequence of scaling factors satisfying $0<\alpha_{1}<\cdots<\alpha_{T_{warm}}<1$. In practice, we empirically find that setting $T_{warm}=4$ provides a stable convergence. We use a predefined sequence of scaling factors: $\alpha=\left\{\frac{1}{100},\frac{1}{50},\frac{1}{10},\frac{1}{5}\right\}$, which produces the epoch-wise regularization strength: $\zeta_{t}\;\in\;\left\{\frac{\zeta}{100},\frac{\zeta}{50},\frac{\zeta}{10},\frac{\zeta}{5},\zeta,\zeta,\dots\right\}$. This warm-up schedule ensures that the regularization penalty is introduced progressively in the initial training epochs without suffering from exploding loss values. Figure 1 shows an overview of the training process with polynomial approximation.
Figure 1: Overview of the two-stage training pipeline of a neural network with polynomial approximation. 1 A ReLU network is approximated using quantization-aware polynomial fitting to obtain an initial polynomial network. 2 The polynomial network is then trained through regularized polynomial network training, consisting of activation regularization, pre-activation clipping, and warm-up scheduling, to ensure stable convergence.
@@ 4.1 Node Fusing
Our first optimization is node fusing, which merges consecutive computational nodes into a single equivalent function. This reduces both the number of operations and multiplicative depth. Formally, two sequential nodes $F(G(x))$ are replaced with $H(x)$ such that $F(G(x))=H(x)$, with $H(x)$ more efficient to evaluate. Following this principle, we present four cases where compositions are fused into a single, more efficient operation, as illustrated in Figure 2. Since all cases involve batch normalization, recall its polynomial form:
$B(x)=b_{1}x+b_{0}$
with $b_{1}=\frac{\gamma}{\sigma},\;b_{0}=\beta_{b}-b_{1}\mu$, where $\mu$, $\sigma$ are the input mean and standard deviation, and $\gamma$, $\beta_{b}$ are learnable parameters. We outline the fusion cases below. See §B.1 for full derivations.
@@
Case 2: $B(C(x))\mapsto C(x)$. Batch normalization can be fused into a preceding convolution [26]. A convolution is given by $C(x)=\sum_{i}w_{i}x_{i}+\beta_{c}$, where $w_{i}$ are the weights and $\beta_{c}$ is the bias. The fused form becomes a single convolution $B(C(x))=\sum_{i}\omega_{i}x_{i}+\alpha$, with fused coefficients $\omega_{i}=b_{1}w_{i},\;\alpha=b_{1}\beta_{c}+b_{0}$.
Case 3: $P(B_{X}(x)+B_{Y}(y))\mapsto S(x,y)$. In residual networks, skip connections often merge two batch-normalized branches by summing them before applying an activation $P(\cdot)$. Specifically, $P(B_{X}(x)+B_{Y}(y))$. Here, $B_{X}$ and $B_{Y}$ are independent batch normalization layers to inputs $x$ and $y$, respectively. This structure can be fused into a single quadratic bivariate polynomial $S(x,y)=d_{X2}x^{2}+d_{Y2}y^{2}+d_{XY}xy+d_{X}x+d_{Y}y+d_{0}$, with coefficients $d_{X2}=c_{2}b_{X1}^{2},\;d_{Y2}=c_{2}b_{Y1}^{2},\;d_{XY}=2c_{2}b_{X1}b_{Y1}$, $d_{X}=b_{X1}(2c_{2}(b_{X0}+b_{Y0})+c_{1}),\;d_{Y}=b_{Y1}(2c_{2}(b_{X0}+b_{Y0})+c_{1})$, $d_{0}=c_{2}(b_{X0}+b_{Y0})^{2}+c_{1}(b_{X0}+b_{Y0})+c_{0}$.
Case 4: $P(B_{X}(x)+y)\mapsto S(x,y)$. Some skip connections use identity shortcuts, where the input $y$ is added directly to the batch-normalized branch $B_{X}(x)$ before applying the activation. This structure can be fused into a quadratic bivariate polynomial $S(x,y)=d_{X2}x^{2}+d_{Y2}y^{2}+d_{XY}xy+d_{X}x+d_{Y}y+d_{0}$, with $d_{X2}=c_{2}b_{X1}^{2},\;d_{Y2}=c_{2},\;d_{XY}=2c_{2}b_{X1}$, $d_{X}=b_{X1}(2c_{2}b_{X0}+c_{1}),\;d_{Y}=2c_{2}b_{X0}+c_{1}$, $d_{0}=c_{2}b_{X0}^{2}+c_{1}b_{X0}+c_{0}$.
Node fusing collapses batch normalizations, activation functions, convolutional layers, and skip connections into single fused polynomials, eliminating redundant nodes and reducing multiplicative depth. This technique is applied after model training but before inference, transforming the model structure into a functionally equivalent, more efficient form.
Figure 2: Illustration of node fusing cases. Nodes are batch normalization ( B ), convolution ( C ), univariate ( P ) and bivariate ( S ) polynomial activations, and addition ( + ).
@@ 4.2 Weight Redistribution
Our second optimization is weight redistribution, a technique that redistributes weights across the network while maintaining functional equivalence. The goal is to reduce the multiplicative depth of selected operations by one. It specifically targets average pooling, batch normalization, and polynomial activation. For average pooling, redistribution eliminates the normalization step by setting the divisor to one. For polynomial functions, coefficients are normalized so that the highest-order coefficient is set to one. For example, the polynomial activation $c_{2}x^{2}+c_{1}x+c_{0}$ is transformed into $x^{2}+\bar{c}_{1}x+\bar{c}_{0}$, reducing its multiplicative depth from two to one, the theoretical minimum for an activation function. These transformations alone break model equivalence. To maintain it, other nodes in the network must compensate for the redistributed weights. We refer to the nodes that initiate redistribution as donors, and to nodes that adjust their parameters to restore equivalence as receivers. We represent the network as a directed graph and traverse it to identify all eligible donor nodes. For each donor, redistribution can be applied either forward or backward, affecting receivers among the donor’s successors or predecessors, respectively, as illustrated in Figure 3.
Update Forward.
@@
Donors. We first consider donor updates in the forward direction. Our goal is to construct a normalized function $\bar{F}(x)$ such that $\upsilon\bar{F}(x)=F(x)$, where $F(x)$ is the original donor function and $\upsilon$ is the update term to be propagated to receivers.
Average Pooling. The average pooling operation is given by $\mu(x)=k^{-1}\sum_{i}{x_{i}}$, where $k$ is the kernel size. In the normalized function we set $k=1$, reducing the operation to $\bar{\mu}(x)=\sum_{i}x_{i}$. Thus, for the equality to be true, we have $\upsilon=k^{-1}$.
@@
Polynomial Functions. Batch normalization and polynomial activation are polynomial functions. For a univariate polynomial $P(x)=\sum_{i=0}^{d}{c_{i}x^{i}}$ of degree $d$, and a normalized polynomial $\bar{P}(x)=x^{d}+\sum_{i=0}^{d-1}{\bar{c}_{i}x^{i}}$, where $\bar{c}_{i}$ are the normalized coefficients, we have $\upsilon=c_{d}$ and $\bar{c}_{i}=c_{i}\upsilon^{-1}$ for $\upsilon\bar{P}(x)=P(x)$ to hold. The extension to the bivariate polynomial activation is analogous and deferred to §B.2.
@@
Receivers. After a donor update, receivers must be adjusted to maintain model equivalence. The original composed function is $G(F(\cdot))$, where $F(\cdot)$ is the donor and $G(\cdot)$ is the receiver. Updating the donor to $\bar{F}(\cdot)$ without compensating the receiver results in $G(\upsilon^{-1}F(\cdot))$, breaking model equivalence. To restore it, we modify the receiver to $\bar{G}(\cdot)$ so that $\bar{G}(\bar{F}(\cdot))=\bar{G}(\upsilon^{-1}F(\cdot))=G(F(\cdot))$. Receivers fall into two categories: kernel functions and polynomial functions.
@@
Kernel Functions. Convolution, linear, and average pooling share the form $K(x)=\sum_{i}w_{i}x_{i}+\beta$ ($w_{i}=k^{-1}$ and $\beta=0$ for average pooling). To maintain model equivalence, we must update the kernel function such that $\bar{K}(\upsilon^{-1}x)=K(x)$ is valid. For that, we set $\bar{w}_{i}=w_{i}\upsilon$ and keep $\bar{\beta}=\beta$.
@@
Polynomial Functions. Batch normalization and polynomial activation update as $\bar{P}(\upsilon^{-1}x)=P(x)$, implying that $\bar{c}_{i}=c_{i}\upsilon^{i}$. For the bivariate polynomial activation, in $\bar{S}(\upsilon^{-1}x,y)=S(x,y)$ the index $i$ corresponds to the exponent of $x$, while in $\bar{S}(x,\upsilon^{-1}y)=S(x,y)$ it corresponds to the exponent of $y$.
Update Backward.
@@
Donors. Backward updates are essentially the inverse of forward updates. For average pooling, the update is identical, since $\bar{\mu}(\upsilon x)=\upsilon\bar{\mu}(x)$, while for polynomial donors, it must hold that $\bar{P}(\upsilon x)=P(x)$; hence, $\upsilon=c_{d}^{1/d}$ and $\bar{c}_{i}=c_{i}\upsilon^{-i}$.
@@
Receivers. To preserve model equivalence, receivers must satisfy $\bar{G}(x)=\upsilon G(x)$. Kernel receivers update as $\bar{K}(x)=\upsilon K(x)$, which implies that $\bar{w}_{i}=w_{i}\upsilon$ and $\bar{\beta}=\beta\upsilon$. And polynomial receivers update according to $\bar{P}(x)=\upsilon P(x)$; hence, $\bar{c}_{i}=c_{i}\upsilon$.
Figure 3: Illustration of weight redistribution cases. $D$ represents a donor and $R$ a receiver. The number of bars ($\bar{\cdot}$) on a $D$ or $R$ represents the number of updates the node received. $\mathbf{f}$ represents a forward and $\mathbf{b}$ a backward update.
Weight redistribution is performed after model training, specifically immediately following node fusing, and prior to inference. It reduces multiplicative depth, while maintaining model equivalence. Figure 3 illustrates common cases for CNNs, including residual connections, where redistribution must be applied consistently along all converging or diverging branches. Refer to §B.2 for detailed derivations for all cases.